It is September 2019 and Sapphire Ltd has just paid a dividend of $1.20 a share. Investors require a 12% p.a. return on Sapphire Ltd shares. What would a share in Sapphire Ltd be expected to sell for today (September, 2019) if the dividend is expected to increase by 20% in September, 2020, 15% in September, 2021, 10% in September, 2022 and thereafter by 5% a year forever from September, 2023 onwards?
Answer #1
Year Dividend PV @ 12% PV of dividends
2020 1.44 0.8929 1.2857
2021 1.656 0.7972 1.3202
2022 1.8216 0.7118 1.2966
Total 3.9024

Price at the end of 2022 = (1.8216* 105%) / (0.12 - 0.05) = 27.324

Present value of the price = 27.324 * 0.7118 = 19.4487

Current stock price = 3.9024 + 19.4487 = 23.3511

A share in Sapphire Ltd be expected to sell for today for $23.3511.

Any doubt please comment.